# R2C2-Coder: Enhancing and Benchmarking Real-world Repository-level Code Completion Abilities of Code Large Language Models

### v1

@@ Abstract
Code completion models have made significant progress in recent years. Recently, repository-level code completion has drawn more attention in modern software development, and several baseline methods and benchmarks have been proposed. However, existing repository-level code completion methods often fall short of fully using the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies. Besides, the existing benchmarks usually focus on limited code completion scenarios, which cannot reflect the repository-level code completion abilities well of existing methods. To address these limitations, we propose the R2C2-Coder to enhance and benchmark the real-world repository-level code completion abilities of code Large Language Models, where the R2C2-Coder includes a code prompt construction method R2C2-Enhance and a well-designed benchmark R2C2-Bench. Specifically, first, in R2C2-Enhance, we first construct the candidate retrieval pool and then assemble the completion prompt by retrieving from the retrieval pool for each completion cursor position. Second, based on R2C2-Enhance, we can construct a more challenging and diverse R2C2-Bench with training, validation and test splits, where a context perturbation strategy is proposed to simulate the real-world repository-level code completion well. Extensive results on multiple benchmarks demonstrate the effectiveness of our R2C2-Coder.
@@ 1 Introduction
Large Language Models for Code (Code LLMs), such as Codex [12], CodeGen [40, 41], and StarCoder [32], have shown their power to enhance developer productivity as their promising results in code completion [4, 5, 12, 39]. Besides, to evaluate these models in real-world scenarios, multiple code completion benchmarks are proposed. Recently, the challenging repository-level code completion has drawn more attention, where the extensive cross-file dependencies (i.e., contextual information from other code files within the same repository) are used to enhance the completion performance. Therefore, a direct but complex solution is to enhance the context length of LLMs [13, 14, 43]. Nonetheless, increasing the context length significantly raises costs and is not affordable, especially concerning the large number of files in a given repository. Hence, several recent baseline methods [16, 37] apply the widely-used Retrieval-Augmented Generation (RAG) strategy by retrieving the most relevant code snippets to improve the repository-level code completion.
However, this vanilla strategy exhibits significant limitations. First, it usually preprocesses the source code files into code snippets under a repository in Fig. 1, and performs retrieval on these code fragments for retrieval query. However, these fragments only preserve local and discrete information and fail to cover more diverse and fundamental contexts of program semantics, such as the intricate network of dependencies, shared utility functions, definition of the input and output parameters, etc. Besides, as retrieved code snippets based on retrievers (e.g., BM25 [48]) usually contain irrelevant or noisy contexts, which may affect the completion results, it is important to evaluate and improve the robustness of the noisy contexts for Code LLMs. Moreover, we observe that the existing repository-level code completion benchmarks focus on limited completion scenes and cannot fully simulate real-world usage well (See Section 4).
@@ 2 Related Works
Code Large Language Models. The successive achievements of generative language modeling [9, 10, 11, 46, 55, 56, 6, 60, 35, 53, 23, 59, 22, 64, 18] have inspired extensive studies on generative AI for software engineering. While closed-source models [1, 12, 15] achieve significant dominance in the benchmark leaderboards [12, 24, 36], their inaccessible model checkpoints and source code are detrimental to subsequent innovation. In contrast, plenty of open-source models, for instance, Code T5 [57, 58], CodeGen [40, 41], StarCoder [3, 32, 38], Code Llama [49] and DeepSeekCoder [21], have been proposed and have substantially driven the development of code intelligence.
Repository-level Code Completion. Existing methods are usually based on a retrieval augmented generation (RAG), and we categorize related works into ranking-based and fusion-based methods. Ranking-based approaches [34, 42, 44, 63] explicitly select those code snippets in the repository that are highly similar to the incomplete code. In contrast, fusion-based methods [17, 50, 51] focus on organizing the repository-level context to be jointly modeled with the language model and allow the model to automatically select the most relevant information.
Code Benchmark. Program synthesis is to prompt code models to solve programming problems using the input description. The generated code snippet should be functional to pass all test cases [4, 5, 12, 19, 31, 36, 62]. Besides, numerous benchmarks can be used to comprehensively evaluate code models, such as code translation [29, 61, 65], code retrieval [25, 27, 33, 39], and vulnerability repair [26, 45, 47, 54]. Recently, several benchmarks have been proposed on repository-level code completion [2, 3, 7, 17, 37, 42, 51, 50, 63]. For example, RepoBench [37] measures the retrieval and completion performance of the system. CrossCodeEval [16] evaluates three cross-file prompting levels: in-file, retrieval, and retrieval within references. However, RepoBench only supports two programming languages and neglects the FIM setting. CrossCodeEval only considers explicit file imports, resulting in few cross-file dependencies and low multi-line prediction difficulty.
@@ 3.1.1 Construction of the Candidate Retrieval Pool
Snippet Context Extraction. For the snippet context, we iteratively scan the files in the repository and extract contiguous $M$ lines (in all our experiments, we set $M=10$ follow [16]) of overlapping code fragments, which are the candidates for context retrieval.
@@ 3.1.2 Construction of the Completion Prompt
Context Retrieval. In Fig. 2, for the current cursor position, we first build the retrieval query based on the prefix and suffix contexts. Specifically, for the current cursor position, we first use the previous $P$ lines and subsequent $S$ lines as the prefix and suffix contexts, respectively. Therefore, the retrieval query usually has $P+S$ lines. Note that if the prefix and suffix have one overlapped line, the retrieval query has $P+S-1$ lines. Then, we use the retriever (e.g., BM 25 [48]) to calculate the similarity between the retrieval query and the candidate retrieval pool. After that, we can select the retrieved contexts based on the similarities. Note that in the retrieval process, we remove the abstract context and snippet context of the current code file, where the cursor position is located.
Prompt Assemble. We assemble the completion prompt by using the current code file, and retrieved contexts. Specifically, we first follow [16] to maintain all context of the current code file as the in-file context, where the completion cursor position is located. Then, we preserve the top-$K$ similar abstract contexts as the coarse-grained global structure information, After that, we append the most relevant snippet context based on the similarity scores until the maximum number of tokens is $N$, where we set $N$ as 4,096 by default. Finally, we can obtain the assembled prompt with $N$ tokens, which will be sent to the Code LLMs for generating the completion results.
@@ 3.2.1 Dataset Generation
Initial Repository Collection. We collect permissively licensed repositories from GitHub. Specifically, following [16], we contain four languages and keep only repositories with the number of stars $>=3$. Then we filter out repositories with fewer than 10 or more than 50 source code files. Finally, we remove these repositories with at least one source code file that exactly matches one of the code files in the Stack [30] dataset. As a result, we ended up with 54972, 51796, 49790, and 35410 repositories for Python, Java, TypeScript, and C#, respectively.
@@ 3.2.2 Dataset Statistics
The statistics of R2C2-Bench are shown in Table 1444We use the StarCoder tokenizer [32] to compute the number of tokens.. We have provided the statistics of CrossCodeEval and validation split of R2C2-Bench in Appendix B.4
@@ 4 Experiments
We conduct experiments on CrossCodeEval [16], CrossCodeEval+, and R2C2-Bench with three popular Code LLMs (i.e., StarCoder-7B [32], DeepSeekCoder-7B [21] and Code Llama-7B [49]) (See Appendix C.1 for more details). For CrossCodeEval+, we collect the original repositories of CrossCodeEval [16] then parse and extract the cross-file dependencies using R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in the original CrossCodeEval (See Appendix C.2 for more details).
@@ 4.1 Experimental Setup
Retrieval in CrossCodeEval. This setup is similar to the “+ Retrieval” setting of CrossCodeEval [16], where the retrieval candidates are constructed by exhaustively scanning files in the same repository and extracting contiguous-$M$-line code fragments ($M=10$). Then, the top 5 similar candidates and the test query assemble the prompt. Note that for better baseline performance, we do not truncate cross-file contents into a total length of 512 tokens as [16].
Retrieval with R2C2-Enhance. Based on R2C2-Enhance in Sec. 3.1, given a repository, we first construct the candidate retrieval pool via abstract and snippet context extraction. Then, we build the retrieval query including prefix length $P=5$ and suffix length $S=5$. After that, candidates are concatenated to the beginning of the in-file context based on their Jaccard similarity [28] in descending order, until the cumulative length, including the in-file context, reaches the maximum token limit of $N=4096$. Note that we set the number of the abstract context $K$ as 3.
@@
Evaluation Metrics. Following CrossCodeEval [16], we evaluate code LLMs with code match metrics including exact match (EM) and edit similarity (ES). Exact match expects that the prediction and the real code are extremely identical, while edit similarity flexibly allows the generated samples to match the original code through a certain number of editing operations.
@@ 4.2 Results
We present results on R2C2-Bench in Table 3.2.2, CrossCodeEval [16] in Table 4.2, and CrossCodeEval+ in Table 4.2. Based on Table 4.2 and Table 4.2, cross-file context is undoubtedly efficacious, while retrieval with R2C2-Enhance surely exhibits its superiority compared with retrieval method in CrossCodeEval. Moreover, fine-tuning strategy significantly improves the performance on both benchmarks. Besides, when abstract contexts are removed from the candidate retrieval pool, these code LLMs suffer from degraded performance. Notably, in Table 3.2.2, after applying R2C2-Enhanced tuning, the Code Llama-7B, which is the worst in the in-file setting, outperforms the StarCoder-7B. Overall, we could say that there exists a performance promotion path, “In-file” $\rightarrow$ “+ R2C2-Enhance w.o. Abs.” $\rightarrow$ “+ R2C2-Enhance” $\rightarrow$ “+ R2C2-Enhanced Tuning”.
@@ 4.3 Analysis
Ablations. We conduct ablation studies on R2C2-Enhance and present the results in Table 4.2. We fine-tune StarCoder-7B [32] with R2C2-Enhance and report the performance on the validation set of R2C2-Bench. Meanwhile, we conduct experiments on the retrieval manner, where we replace the FIM mode with “prefix-only”. We could find that suffix contexts are necessary as the results of $P=5,S=3$ even surpass those of $P=10$. Additionally, we remove either snippet contexts (“w.o. Snippet Contexts”) or abstract contexts (“w.o. Abstract Contexts”) simultaneously in fine-tuning and testing. Results on various prefix lengths ($P$) and suffix lengths ($S$) are further reported for the ablation model. Despite the larger impact without using snippet contexts, abstract contexts still deliver 3.5% and 3.4% performance gains. Furthermore, a case study in Fig. 6 shows how the code LLMs benefit from abstract context extraction.
Context Perturbation Helps Cope with Confusing Contexts. In real-world scenarios, retrievers would produce low-quality results as shown in Fig. 7. To enhance the model robustness on irrelevant contexts, we select $Q\%$ of the cursors and randomly remove the top similar contexts during R2C2-Enhanced tuning. Fig. 4 visualizes how the performance of StarCoder-7B on EM and ES scales when using different perturbation rates for R2C2-Enhanced tuning. The perturbation rate $Q\%$ for fine-tuning varies in $0\%,10\%,20\%,50\%,$ and $80\%$, while the perturbation rate for the validation set is always kept as 10%. The best results are achieved under perturbation rate $Q\%=10\%$, which is also the default setting of R2C2-Enhanced tuning for Table 3.2.2 , 4.2 and 4.2.
Similarity Metric Python Java TypeScript C# Average Execution Time (ms)
@@
Lexical Retrievers are Cost-effective. For ranking-based code retrieval, several approaches could be adopted as the similarity metric. We categorize the metrics into lexical ones, including Jaccard similarity [28] and BM25 [48], and neural ones such as UniXcoder [20]. In real-world scenarios, we often need to make a trade-off, as lexical retrievers could save more response time (RT) while neural retrievers would theoretically achieve higher performance. To find the most suitable retriever for R2C2-Enhance, we evaluate the code match performance and retrieval duration with Jaccard similarity, BM25, and UniXCoder. Table 4.3 presents the code match performance and retrieval duration of each retriever. Although UniXCoder outperforms BM25 on code match, its time consumption is extremely worse than lexical retrievers. We find that Jaccard similarity is the best retriever regarding both effectiveness and efficiency. Thus, we choose Jaccard similarity by default. See Appendix C.3 for details of execution time benchmarking.
Scalability of Model Performance. We present how the performance of StarCoder scales w.r.t model parameters in Table 4.3. We find that almost no performance gain is obtained for StarCoder-3B when using R2C2-Enhance with abstract and snippet contexts. After R2C2-Enhanced tuning, StarCoder-3B can outperform the inference-only results of StarCoder-7B. In Table 4.3 we further evaluate the fine-tuned StarCoder-7B when using different sizes of tuning datasets based on our R2C2-Enhanced tuning. We observe that when increasing from 10K samples to 100K samples per language, better results are obtained. When continually increasing, the results are relatively stable. Thus, we choose 10K samples per language as the default training set size of R2C2-Bench.
@@
Limitations of CrossCodeEval. We summarize three major drawbacks of CrossCodeEval [16] as follows: 1) The retrieval results provided in CrossCodeEval are produced by sliding-window-based searching, which could potentially miss long-term dependencies that exceed the length of $M$. In contrast, abstract extraction expands the reception field of code LLMs with coarse-grained global information. 2) CrossCodeEval only considers explicit dependencies (i.e., file imports), but there could be implicit references during software development. For example, in Java, usages in the same package do not require the import statement. As shown in Fig. 8, such implicit dependencies could not be detected by Ding et al. [16], but are successfully covered by our R2C2-Enhance during constructing CrossCodeEval+. 3) Compared with CrossCodeEval, our R2C2-Bench provides more complex cross-file dependencies (see “Avg. # cross files in prompt” in Table 1 and Table B.4) and more difficult multi-line test samples in Fig. 5).
@@
Training samples (×$\times 4$) Python Java TypeScript C# Average

### v3

@@ Abstract
Recently, repository-level code completion has drawn more attention in modern software development, and several baseline methods and benchmarks have been proposed. However, existing repository-level code completion methods often fall short of fully using the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies. Besides, the existing benchmarks usually focus on limited code completion scenarios, which cannot reflect the repository-level code completion abilities well of existing methods. To address these limitations, we propose the R2C2-Coder to enhance and benchmark the real-world repository-level code completion abilities of code Large Language Models, where the R2C2-Coder includes a code prompt construction method R2C2-Enhance and a well-designed benchmark R2C2-Bench. Specifically, first, in R2C2-Enhance, we first construct the candidate retrieval pool and then assemble the completion prompt by retrieving from the retrieval pool for each completion cursor position. Second, based on R2C2-Enhance, we can construct a more challenging and diverse R2C2-Bench with training, validation and test splits, where a context perturbation strategy is proposed to simulate the real-world repository-level code completion well. Extensive results on multiple benchmarks demonstrate the effectiveness of our R2C2-Coder.
@@ 1 Introduction
Large Language Models for Code (Code LLMs), such as Codex (Chen et al., 2021), CodeGen (Nijkamp et al., 2023a, b), and StarCoder (Li et al., 2023), have shown their power to enhance developer productivity as their promising results in code completion (Athiwaratkun et al., 2023; Austin et al., 2021; Chen et al., 2021; Lu et al., 2021). Besides, to evaluate these models in real-world scenarios, multiple code completion benchmarks are proposed. Recently, the challenging repository-level code completion has drawn more attention, where the extensive cross-file dependencies (i.e., contextual information from other code files within the same repository) are used to enhance the completion performance. Therefore, a direct but complex solution is to enhance the context length of LLMs (Chen et al., 2023a, b; Peng et al., 2023). Nonetheless, increasing the context length significantly raises costs and is not affordable, especially concerning the large number of files in a given repository. Hence, several recent baseline methods (Ding et al., 2023; Liu et al., 2023b) apply the widely-used Retrieval-Augmented Generation (RAG) strategy by retrieving the most relevant code snippets to improve the repository-level code completion.
However, this vanilla strategy exhibits significant limitations. First, it usually preprocesses the source code files into code snippets under a repository in Fig. 1, and performs retrieval on these code fragments for retrieval query. However, these fragments only preserve local and discrete information and fail to cover more diverse and fundamental contexts of program semantics, such as the intricate network of dependencies, shared utility functions, definition of the input and output parameters, etc. Besides, as retrieved code snippets based on retrievers (e.g., BM25 Robertson et al. (2009)) usually contain irrelevant or noisy contexts, which may affect the completion results, it is important to evaluate and improve the robustness of the noisy contexts for Code LLMs. Moreover, we observe that the existing repository-level code completion benchmarks focus on limited completion scenes and cannot fully simulate real-world usage well (See Section 4).
@@ 2 Related Works
Code Large Language Models. The successive achievements of generative language modeling (Black et al., 2022, 2021; Brown et al., 2020; Radford et al., 2019; Touvron et al., 2023; Wang and Komatsuzaki, 2021; Bai et al., 2024; Wu et al., 2024; Liu et al., 2024; Sun et al., 2024; Guo et al., 2024b; Wang et al., 2023b; Guo et al., 2023; Zhang et al., 2024; Du et al., 2024) have inspired extensive studies on generative AI for software engineering. While closed-source models (Achiam et al., 2023; Chen et al., 2021; Chowdhery et al., 2023) achieve significant dominance in the benchmark leaderboards (Chen et al., 2021; Hendrycks et al., 2020; Liu et al., 2023a), their inaccessible model checkpoints and source code are detrimental to subsequent innovation. In contrast, plenty of open-source models, for instance, Code T5 (Wang et al., 2023a, 2021), CodeGen (Nijkamp et al., 2023a, b), StarCoder (Allal et al., 2023; Li et al., 2023; Lozhkov et al., 2024), Code Llama (Roziere et al., 2023) and DeepSeekCoder (Guo et al., 2024a), have been proposed and have substantially driven the development of code intelligence.
Repository-level Code Completion. Existing methods are usually based on a retrieval augmented generation (RAG), and we categorize related works into ranking-based and fusion-based methods. Ranking-based approaches (Liao et al., 2023; Pei et al., 2023; Phan et al., 2024; Zhang et al., 2023) explicitly select those code snippets in the repository that are highly similar to the incomplete code. In contrast, fusion-based methods (Ding et al., 2022; Shrivastava et al., 2023a, b) focus on organizing the repository-level context to be jointly modeled with the language model and allow the model to automatically select the most relevant information.
Code Benchmark. Program synthesis is to prompt code models to solve programming problems using the input description. The generated code snippet should be functional to pass all test cases (Athiwaratkun et al., 2023; Austin et al., 2021; Chen et al., 2021; Gu et al., 2024; Lai et al., 2023; Liu et al., 2023a; Yu et al., 2024). Besides, numerous benchmarks can be used to comprehensively evaluate code models, such as code translation (Jiao et al., 2023; Yan et al., 2023; Zhu et al., 2022), code retrieval (Huang et al., 2021; Husain et al., 2019; Li et al., 2024; Lu et al., 2021), and vulnerability repair (Huq et al., 2022; Prenner and Robbes, 2023; Richter and Wehrheim, 2022; Tian et al., 2024). Recently, several benchmarks have been proposed on repository-level code completion (Agrawal et al., 2023; Allal et al., 2023; Bairi et al., 2023; Ding et al., 2022; Liu et al., 2023b; Pei et al., 2023; Shrivastava et al., 2023b, a; Zhang et al., 2023). For example, RepoBench (Liu et al., 2023b) measures the retrieval and completion performance of the system. CrossCodeEval (Ding et al., 2023) evaluates three cross-file prompting levels: in-file, retrieval, and retrieval within references. However, RepoBench only supports two programming languages and neglects the FIM setting. CrossCodeEval only considers explicit file imports, resulting in few cross-file dependencies and low multi-line prediction difficulty.
@@ 3.1.1 Construction of the Candidate Retrieval Pool
Snippet Context Extraction. For the snippet context, we iteratively scan the files in the repository and extract contiguous $M$ lines (in all our experiments, we set $M=10$ follow Ding et al. (2023)) of overlapping code fragments, which are the candidates for context retrieval.
@@ 3.1.2 Construction of the Completion Prompt
Context Retrieval. In Fig. 2, for the current cursor position, we first build the retrieval query based on the prefix and suffix contexts. Specifically, for the current cursor position, we first use the previous $P$ lines and subsequent $S$ lines as the prefix and suffix contexts, respectively. Therefore, the retrieval query usually has $P+S$ lines. Note that if the prefix and suffix have one overlapped line, the retrieval query has P+S1P+S-1 lines. Then, we use the retriever (e.g., BM 25 (Robertson et al., 2009)) to calculate the similarity between the retrieval query and the candidate retrieval pool. After that, we can select the retrieved contexts based on the similarities. Note that in the retrieval process, we remove the abstract context and snippet context of the current code file, where the cursor position is located.
Prompt Assemble. We assemble the completion prompt by using the current code file, and retrieved contexts. Specifically, we first follow Ding et al. (2023) to maintain all context of the current code file as the in-file context, where the completion cursor position is located. Then, we preserve the top-$K$ similar abstract contexts as the coarse-grained global structure information, After that, we append the most relevant snippet context based on the similarity scores until the maximum number of tokens is $N$, where we set $N$ as 4,096 by default. Finally, we can obtain the assembled prompt with $N$ tokens, which will be sent to the Code LLMs for generating the completion results.
@@ 3.2.1 Dataset Generation
Initial Repository Collection. We collect permissively licensed repositories from GitHub. Specifically, following (Ding et al., 2023), we contain four languages and keep only repositories with the number of stars >=3>=3. Meanwhile, we exclude repositories that contain fewer than 10 or more than 50 source code files, and we eliminate any repositories that have at least one source code file identical to any file found in the Stack dataset (Kocetkov et al., 2022). As a result, we ended up with 54972, 51796, 49790, and 35410 repositories for Python, Java, TypeScript, and C#, respectively.
@@ 3.2.2 Dataset Statistics
The statistics of R2C2-Bench are shown in Table 1444We use the StarCoder tokenizer (Li et al., 2023) to compute the number of tokens.. We have provided the statistics of CrossCodeEval and validation split of R2C2-Bench in Appendix B.4
@@ 4 Experiments
We conduct experiments on CrossCodeEval (Ding et al., 2023), CrossCodeEval+, and R2C2-Bench with three popular Code LLMs (i.e., StarCoder-7B (Li et al., 2023), DeepSeekCoder-7B (Guo et al., 2024a) and Code Llama-7B (Roziere et al., 2023)) (See Appendix C.1 for more details). For CrossCodeEval+, we collect the original repositories of CrossCodeEval (Ding et al., 2023) then parse and extract the cross-file dependencies using R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in the original CrossCodeEval (See Appendix C.2 for more details).
@@ 4.1 Experimental Setup
Retrieval in CrossCodeEval. This setup is similar to the “+ Retrieval” setting of CrossCodeEval (Ding et al., 2023), where the retrieval candidates are constructed by exhaustively scanning files in the same repository and extracting contiguous-$M$-line code fragments ($M=10$). Then, the top 5 similar candidates and the test query assemble the prompt. Note that for better baseline performance, we do not truncate cross-file contents into a total length of 512 tokens as (Ding et al., 2023).
Retrieval with R2C2-Enhance. Based on R2C2-Enhance in Sec. 3.1, given a repository, we first construct the candidate retrieval pool via abstract and snippet context extraction. Then, we build the retrieval query including prefix length $P=5$ and suffix length $S=5$. After that, candidates are concatenated to the beginning of the in-file context based on their Jaccard similarity (Jaccard, 1912) in descending order, until the cumulative length, including the in-file context, reaches the maximum token limit of $N=4096$. Note that we set the number of the abstract context $K$ as 3.
@@
Evaluation Metrics. Following CrossCodeEval (Ding et al., 2023), we evaluate code LLMs with code match metrics including exact match (EM) and edit similarity (ES). Exact match expects that the prediction and the real code are extremely identical, while edit similarity flexibly allows the generated samples to match the original code through a certain number of editing operations.
@@ 4.2 Results
We present results on R2C2-Bench in Table 2, CrossCodeEval (Ding et al., 2023) in Table 3, and CrossCodeEval+ in Table 4. Based on Table 3 and Table 4, cross-file context is undoubtedly efficacious, while retrieval with R2C2-Enhance surely exhibits its superiority compared with retrieval method in CrossCodeEval. Moreover, fine-tuning strategy significantly improves the performance on both benchmarks. Besides, when abstract contexts are removed from the candidate retrieval pool, these code LLMs suffer from degraded performance. Notably, in Table 2, after applying R2C2-Enhanced tuning, the Code Llama-7B, which is the worst in the in-file setting, outperforms the StarCoder-7B. Overall, we could say that there exists a performance promotion path, “In-file” $\rightarrow$ “+ R2C2-Enhance w.o. Abs.” $\rightarrow$ “+ R2C2-Enhance” $\rightarrow$ “+ R2C2-Enhanced Tuning”.
@@ 4.3 Analysis
Ablations. We conduct ablation studies on R2C2-Enhance and present the results in Table 5. We fine-tune StarCoder-7B (Li et al., 2023) with R2C2-Enhance and report the performance on the validation set of R2C2-Bench. Meanwhile, we conduct experiments on the retrieval manner, where we replace the FIM mode with “prefix-only”. We could find that suffix contexts are necessary as the results of $P=5,S=3$ even surpass those of $P=10$. Additionally, we remove either snippet contexts (“w.o. Snippet Contexts”) or abstract contexts (“w.o. Abstract Contexts”) simultaneously in fine-tuning and testing. Results on various prefix lengths ($P$) and suffix lengths ($S$) are further reported for the ablation model. Despite the larger impact without using snippet contexts, abstract contexts still deliver 3.5% and 3.4% performance gains. Furthermore, a case study in Fig. 6 shows how the code LLMs benefit from abstract context extraction.
Context Perturbation Helps Cope with Confusing Contexts. In real-world scenarios, retrievers would produce low-quality results as shown in Fig. 7. To enhance the model robustness on irrelevant contexts, we select $Q\%$ of the cursors and randomly remove the top similar contexts during R2C2-Enhanced tuning. Fig. 4 visualizes how the performance of StarCoder-7B on EM and ES scales when using different perturbation rates for R2C2-Enhanced tuning. The perturbation rate $Q\%$ for fine-tuning varies in $0\%,10\%,20\%,50\%,$ and $80\%$, while the perturbation rate for the validation set is always kept as 10%. The best results are achieved under perturbation rate $Q\%=10\%$, which is also the default setting of R2C2-Enhanced tuning for Table 2 , 3 and 4.
Similarity Metric Python Java TypeScript C# Average
Execution Time
@@
Lexical Retrievers are Cost-effective. For ranking-based code retrieval, several approaches could be adopted as the similarity metric. We categorize the metrics into lexical ones, including Jaccard similarity (Jaccard, 1912) and BM25 (Robertson et al., 2009), and neural ones such as UniXcoder (Guo et al., 2022). In real-world scenarios, we often need to make a trade-off, as lexical retrievers could save more response time (RT) while neural retrievers would theoretically achieve higher performance. To find the most suitable retriever for R2C2-Enhance, we evaluate the code match performance and retrieval duration with Jaccard similarity, BM25, and UniXCoder. Table 6 presents the code match performance and retrieval duration of each retriever. Although UniXCoder outperforms BM25 on code match, its time consumption is extremely worse than lexical retrievers. We find that Jaccard similarity is the best retriever regarding both effectiveness and efficiency. Thus, we choose Jaccard similarity by default. See Appendix C.4 for details of execution time benchmarking.
Scalability of Model Performance. We present how the performance of StarCoder scales w.r.t model parameters in Table 7. We find that almost no performance gain is obtained for StarCoder-3B when using R2C2-Enhance with abstract and snippet contexts. After R2C2-Enhanced tuning, StarCoder-3B can outperform the inference-only results of StarCoder-7B. In Table 8 we further evaluate the fine-tuned StarCoder-7B when using different sizes of tuning datasets based on our R2C2-Enhanced tuning. We observe that when increasing from 10K samples to 100K samples per language, better results are obtained. When continually increasing, the results are relatively stable. Thus, we choose 10K samples per language as the default training set size of R2C2-Bench.
@@
Limitations of CrossCodeEval. We summarize three major drawbacks of CrossCodeEval (Ding et al., 2023) as follows: 1) The retrieval results provided in CrossCodeEval are produced by sliding-window-based searching, which could potentially miss long-term dependencies that exceed the length of $M$. In contrast, abstract extraction expands the reception field of code LLMs with coarse-grained global information. 2) CrossCodeEval only considers explicit dependencies (i.e., file imports), but there could be implicit references during software development. For example, in Java, usages in the same package do not require the import statement. As shown in Fig. 8, such implicit dependencies could not be detected by Ding et al. (2023), but are successfully covered by our R2C2-Enhance during constructing CrossCodeEval+. 3) Compared with CrossCodeEval, our R2C2-Bench provides more complex cross-file dependencies (see “Avg. # cross files in prompt” in Table 1 and Table B.4) and more difficult multi-line test samples in Fig. 5).
@@
Training samples
($\times 4$)
Python Java TypeScript C# Average